cos^3a.cos 3a+sin^3a.sin3a=cos^3 2a
1
Expert's answer
2019-05-07T04:27:17-0400

You need the following identities: 


"1. cos^2 X + sin^2 X = 1""2. sin(X+Y) = sin X cosY + sin Y cos X""3. cos(X-Y) = cos X cos Y + sin X sin Y""4. sin(2X)= 2*sin X cos X"


Solution:


"cos^3 (A) cos (3A) + sin^3 (A) sin (3A) =""cos^2 (A) (cos A cos (3A)) + sin^2 (A)(sin A sin (3A)) ="


Formula 1 gives:


"cos^2 (A) = 1 - sin^2 (A)"


and


"sin^2 (A) = 1 - cos^2 (A)"


"(1-sin^2 (A)) (cos (A) cos (3A)) + (1-cos^2 (A))(sin (A) sin (3A)) =""cos (A) cos (3A) + sin (A) sin (3A) - sin^2 (A) cos (A) cos (3A) - cos^2 (A) sin (A) sin (3A) =""cos (A) cos (3A) + sin (A) sin( 3A) - sin( A )cos (A)(sin (A) cos (3A) + sin (3A) cos(A))="

Formua 2 gives:


"sin (A) cos (3A) + sin (3A) cos (A) = sin (4A)"

Formua 3 gives:


"cos (A) cos (3A) + sin (A) sin (3A) = cos (2A)""cos (2A) - sin (A) cos (A) (sin (4A)) ="


Formua 4 gives:


"sin (4A) = 2*sin (2A) cos (2A)""2*sin (A) cos (A) = sin (2A)""cos (2A) - 2*sin (A) cos (A) (sin (2A) cos (2A)) =""cos (2A) - sin (2A) sin (2A) cos( 2A) =""cos (2A) (1 - sin^2 (2A)) ="


Using Formua 1 again :


"cos( 2A) (cos^2 (2A)) ="


"cos^3 (2A)"